Joe traveled against the wind in a small plane for 3 hr. The return trip with the the wind took 2.8 hr. Find the speed of the wind if the speed of the plane is still air is 180 mph.
Answer by mananth(16949)

Joe traveled against the wind in a small plane for 3 hr. The return trip with the the wind took 2.8 hr. Find the speed of the wind if the speed of the plane is still air is 180 mph.
let the speed of the wind be x mph
Against wind the speed will be 180-x
distance = speed * time
= 3(180-x)
With wind the speed will be x+180
distance = 2.8(x+180)
3(180-x)= 2.8(x+180)
540-3x=2.8x+504
5.8x=36
x=36/5.8
speed of wind = 6.2 mph